OpenStudy (anonymous):

Simplify: (x +2)(x^2 +2x + 3)

OpenStudy (anonymous):

do you know what the distributive property is?

OpenStudy (anonymous):

@DJ29 are you there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you have to distribute the first set of parentheses into the second.

OpenStudy (anonymous):

x*x^2=x^3 x*2x=2x^2 x*3=3x 2*x^2=2x^2 2*2x=4x 2*3=6 now combine like terms

OpenStudy (anonymous):

@DJ29 do you know how to combine like terms?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

combine the like terms from above and thats your answer

OpenStudy (anonymous):

is it x^3 + 4x^2 + 7x + 6

OpenStudy (anonymous):

yup is that one of your choices?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats it! Hope this helped and good luck